OpenStudy (anonymous):

Find the solution of the square root of the quantity of x plus 2 plus 4 equals 6 and determine if it is an extraneous solution. x = 2; extraneous x = 2; not extraneous x = 30; extraneous x = 30; not extraneous

OpenStudy (lynfran):

\[\sqrt{x+2+4}=6\]\[(\sqrt{x+6})^{2}=6^{2}\]\[x+6=36\]\[x=36-6\]

OpenStudy (anonymous):

so x=30; Non-extraneous?

OpenStudy (lynfran):

correct